OpenStudy (moose213):

4 questions please help will fan and medal! 1. Solve the equation using the distributive property and properties of equality. 1/2 (x+6) = 18 What is the value of x? 2. What is the value of x in the equation below? 6 (x-8) = 72 3. Solve the equation using the distributive property and properties of equality. -5 (a+3) = -55 What is the value of a? 4. Consider the solution to the linear equation. 5 (x+6) = 50 5x +30 = 50 5x = 20 x=4 Which describes the inverse operations used after the distributive property?

OpenStudy (moose213):

@Qwertty123

OpenStudy (moose213):

@jjcoolja12

OpenStudy (moose213):

@Tennis5518

OpenStudy (moose213):

@ash2326

OpenStudy (moose213):

Please help =)

OpenStudy (jjcoolja12):

1 Let's solve your equation step-by-step. 1 2 (x+6)=18 Step 1: Simplify both sides of the equation. 1 2 (x+6)=18 Simplify: (Show steps) 1 2 x+3=18 Step 2: Subtract 3 from both sides. 1 2 x+3−3=18−3 1 2 x=15 Step 3: Multiply both sides by 2. 2*( 1 2 x)=(2)*(15) x=30 Answer: x=30

OpenStudy (moose213):

#2?

OpenStudy (moose213):

@jjcoolja12 it would mean the world if you can answer these 4, you already answered 1 =)

OpenStudy (jjcoolja12):

#2 Let's solve your equation step-by-step. 6(x−8)=72 Step 1: Simplify both sides of the equation. 6(x−8)=72 Simplify: (Show steps) 6x−48=72 Step 2: Add 48 to both sides. 6x−48+48=72+48 6x=120 Step 3: Divide both sides by 6. 6x 6 = 120 6 x=20

OpenStudy (moose213):

#3?

OpenStudy (jjcoolja12):

Let's solve your equation step-by-step. −5(a+3)=−55 Step 1: Simplify both sides of the equation. −5(a+3)=−55 Simplify: (Show steps) −5a−15=−55 Step 2: Add 15 to both sides. −5a−15+15=−55+15 −5a=−40 Step 3: Divide both sides by -5. −5a −5 = −40 −5 a=8

OpenStudy (moose213):

the fourth one?

OpenStudy (jjcoolja12):

Let's solve your equation step-by-step. 5(x+6)=50 Step 1: Simplify both sides of the equation. 5(x+6)=50 Simplify: (Show steps) 5x+30=50 Step 2: Subtract 30 from both sides. 5x+30−30=50−30 5x=20 Step 3: Divide both sides by 5. 5x 5 = 20 5 x=4 im not sure what 4 means

OpenStudy (moose213):

addition then division subtraction then division multiplication then subtraction division then addition

OpenStudy (moose213):

Which one for number 4

OpenStudy (moose213):

Which describes the inverse operations used after the distributive property?

OpenStudy (jjcoolja12):

divide

OpenStudy (moose213):

Please help with this last one. I will do ANYTHING, pleaseeee.

OpenStudy (moose213):

So division then addition?

OpenStudy (moose213):

or addition then division?

OpenStudy (moose213):

Or subtraction then division?

OpenStudy (jjcoolja12):

i believe so subtraction then division

OpenStudy (moose213):

One more pweasseeee.

OpenStudy (moose213):

7 (x-3) = 28 What is the value of x?

OpenStudy (moose213):

@jjcoolja12

OpenStudy (jjcoolja12):

Let's solve your equation step-by-step. 7(x−3)=28 Step 1: Simplify both sides of the equation. 7(x−3)=28 Simplify: (Show steps) 7x−21=28 Step 2: Add 21 to both sides. 7x−21+21=28+21 7x=49 Step 3: Divide both sides by 7. 7x 7 = 49 7 x=7 Answer: x=7